OpenStudy (anonymous):

Which of the following is a function equivalent to f(x) = 4(1.8)2x?

OpenStudy (anonymous):

f(x) = 4(3.24)x f(x) = 4(3.24x) f(x) = 12.96x f(x) = 51.84x

OpenStudy (anonymous):

@Vocaloid @pooja195

Vocaloid (vocaloid):

f(x) = 4(1.8)^2(x) keep the 4 alone, since its not attached to the exponent (1.8)^2 = ?

OpenStudy (anonymous):

3.24

OpenStudy (anonymous):

then its 3.24*4 right?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

no, you can't multiply the 4 because 3.24 is attached to the exponent x your answer will be f(x) = 4(3.24)^x

OpenStudy (anonymous):

oh ok thank u and btw do u have instagram

Vocaloid (vocaloid):

no